# From tests to effect sizes: Quantifying uncertainty and statistical variability in multilingual and multitask NLP evaluation benchmarks

### v1

@@ Abstract
In this paper, we introduce a set of resampling-based methods for quantifying uncertainty and statistical precision of evaluation metrics in multilingual and/or multitask NLP benchmarks. We show how experimental variation in performance scores arises from both model- and data-related sources, and that accounting for both of them is necessary to avoid substantially underestimating the overall variability over hypothetical replications. Using multilingual question answering, machine translation, and named entity recognition as example tasks, we also demonstrate how resampling methods are useful for computing sampling distributions for various quantities used in leaderboards such as the average/median, pairwise differences between models, and rankings.
From tests to effect sizes: Quantifying uncertainty and statistical variability in multilingual and multitask NLP evaluation benchmarks
@@ 1 Introduction
Over the last several years, multilingual research has undergone exponential growth within NLP, both in terms of model capabilities as well as evaluation datasets. Since no paper can evaluate on all languages, it is important to determine to what extent research findings generalize from one language and from one task to another. This is particularly true in recent years with the broad applicability of large language models (LLMs), where the goal often is to draw conclusions about the models that are not task- or language-dependent.
Despite this need, evaluation setups in multilingual NLP research tend to focus on within-language evaluation and fall into two broad categories. In the first category, the same experiments are repeated across several languages and the results analyzed separately. The inferences are typically reported as one collection, e.g. “Model A significantly outperforms model B on English–Finnish and English–Turkish but not English–German” or “Model A outperforms model B on 13 out of 43 language pairs.” The second variant, depicted in Figure 1, is the “leaderboard-type” scenario in which performance is evaluated on each of the $L$ languages after which the numbers are distilled into a single performance measure using a pre-specified aggregation function such as the arithmetic mean.
Given the aggregate scores, the main interest may then be to rank the $M$ models based on the single score or two compare differences between observed scores. The most interesting questions to practitioners often revolve around whether a given model “truly” outperforms another model or whether an observed performance difference was simply random noise that could be expected to occur under replication when comparing two equally performing models.
@@
From a statistical perspective, estimating this “consistency with random noise” is very similar to determining whether an estimate is statistically significant. Indeed, being 1.96 ($\approx 2$) standard errors from zero is often taken as a rough equivalent to statistical significance (Gelman, 2018). To estimate such an “effect size” requires specifying the source(s) of random variation in our experiments. Typically, this variation has been treated as arising from one of two sources: on the one hand, there is data-side variability which quantifies how differently each model would perform if a slightly different test set were used for a given task. In practice this is typically done using resampled versions of the original test set, for example using the bootstrap resampling algorithm (Efron, 1979).
@@
From a sampling perspective, data-side variability can be interpreted as sampling error associated with the process of constructing a test set for a given task by sampling data from a larger set of all possible test sets. Another source of data-related variability that appears in leaderboards and other multi-task evaluations is the choice of evaluation tasks. When ranking is performed based on aggregated quantities, the overall performance and rankings of models will vary as a function of which tasks are chosen. This can be understood as the sampling error associated with creating a evaluation benchmark out of the larger population of all possible ones that could have been constructed.
@@
On the other hand, there is also model-side variability that will be present even if the data set is held constant. With LLMs, decoding often involves random sampling of tokens, especially with generative tasks like question answering. By drawing multiple responses for each question, the performance will fluctuate randomly around some average. Even if the inference algorithm is deterministic (e.g. beam search), any potential finetuning will likely be nondeterministic due to, for example, random shuffling of minibatches and weight initialization. This will yield slightly different parameters from run to run, which will potentially yield different responses and different values of the performance metric. From a sampling perspective, model-side uncertainty can be seen as drawing samples from the probability distribution over possible responses defined by the model.
@@
In this paper, we provide a new perspective on how resampling-based methods can be useful in analysis of experimental results in multilingual NLP. Our specific contributions as follows.
@@
First, through connections to statistics, we show how experimental variance can be decomposed into between- and within-language components and how the latter further decomposes into model-based and data-based components ($\sigma_{l}$ and $\tau_{l}$) which arise naturally from the structure of the experimental data and can be used in conjunction with each other when replicating experiments.
@@
Second, using QA, machine translation, and NER as case studies, our results demonstrate that none of these sources of variation are negligible. This suggests that only analyzing one source of error may underestimate the total variation and expose researchers to the risk of drawing incorrect conclusions about the relative merits of models. We show that between-language variance accounts for most of the variance in leaderboard-style aggregate evaluations, which suggests that downstream use of an LLM may be better modeled by resampling tasks as well as replications.
@@
Third, we show how resampling and estimates of between and within-language variation can be used to derive uncertainty-aware estimates of “overall performance,” as well as more complex quantities such as relative rankings of models and approximate distributions of pairwise performance differences between them. All of our methods run in seconds to minutes and present no major computational bottlenecks on top of the overall inference time complexity.
@@
We provide an implementation of our approach in a toolkit which we make freely available at www.github.com/j0ma/reuben. [^1] The name reuben is an acronym for “REsampling-based Uncertainty Bounds for Evaluating NLP.
@@ 2 Related work
In contrast, Ulmer et al. (2022) advocate for understanding model performance variation based on multiple model training runs instead of resampling, and using this model-side variation to assess statistical significance. There is no clear consensus, though, and e.g. Bethard (2022) expressly warn against using random seeds to provide estimates of score distributions based on the argument that random seeds should be optimized like other hyperparameters. Some work in NLP has also evaluated different sources of variation in evaluation of MT metrics (Xiang et al., 2022).
@@
There has also been a healthy debate regarding best practices of test set construction and evaluation using held-out datasets in more generally. Gorman and Bedrick (2019) and others (e.g. Kodner et al., 2023; Liu and Dorr, 2024) argue that we should randomize our splits and evaluate on several versions of our test sets to reduce the chance of false positives. Søgaard et al. (2021) instead favor using approaches based on adversarial splitting, arguing that naive randomization will underestimate the variation.
Benchmarks have also received lots of attention in the general machine learning community. An excellent survey is Dehghani et al. (2021) where the authors discuss the lifecycle of overall ML benchmarks, how they become stale as well as what is random when models are evaluated. The question of what aggregation function to use in leaderboards is also explored by Tatiana and Valentin (2021) where the authors explore using both the arithmetic, geometric and harmonic means. In particular, they show that results may be wildly different depending on what aggregation measure is used. Longjohn et al. (2025) also survey and develop resampling-based and Bayesian methods for popular computer vision benchmarks for LLMs.
@@
There exists a substantial literature (e.g. Nikitin et al., 2024; Da et al., 2025; Chen et al., 2024, 2025; Yang et al., 2024; Ye et al., 2024; Wagner et al., 2024; Blackwell et al., 2024) on a task referred to as uncertainty quantification in NLP, where the task is to estimate some properties of an LLMs distribution over continuations, $p(x|x_{\text{prompt}})$, that predicts whether the model is able to produce the correct the answer to this question. While this task quantifies per-example uncertainty, our efforts instead focus on experimental variation. i.e. estimating the noise we would expect to see under repeated experiments and hypothetical evaluation suites.
@@ 3 Statistical background
Under infinite replications, our observed values $x^{(m)}_{l}$ values will center around some mean value and exhibit a degree of fluctuation around it, representing measurement error arising from both model-side randomness (e.g. nondeterministic decoding) as well as sampling error of the test set. We can represent this as $x^{(m)}_{lr}=\mu^{(m)}_{l}+\varepsilon^{\text{repl}}_{lr}$ where $\varepsilon^{\text{repl}}_{lr}$ represents the deviation of observed scores from the language-specific mean $\mu^{(m)}_{l}$. In addition to this “within-language” variation arising from replication, there is also “between-language” variability related to what tasks are included in the leaderboard. Concretely, each $\mu^{(m)}_{l}$ also differs from the global mean performance across all tasks by some amount, i.e. $\mu^{(m)}_{l}=\bar{\mu}^{(m)}+\varepsilon^{\text{lang}}_{l}$. This lets us decompose the overall observations into $x^{(m)}_{lr}=\bar{\mu}^{(m)}+\varepsilon^{\text{lang}}_{l}+\varepsilon^{\text{repl}}_{lr}$. Taking the variance of both sides, we obtain a first-principles variance decomposition into between- and within-language components $\mathbb{V}[x^{(m)}_{lr}]=\mathbb{V}[\varepsilon^{\text{lang}}_{l}]+\mathbb{V}[\varepsilon^{\text{repl}}_{lr}]=\nu_{m}^{2}+\eta^{2}_{lm}$ where $\nu_{m}^{2}$ and $\eta^{2}_{lm}$ refer to between- and within-language variation of model $m$, respectively.
@@
In our analysis, we use both seed- and bootstrap-based resampling to yield a total of $R=SB$ replications per language. While the $B$ resamples may be much cheaper to obtain than the $S$ model re-instantiations in terms of time cost, we feel that using both is helpful to properly estimate the decomposition as well as to understand whether the sources contribute equally to the total within-language variance for each language.
@@
Most immediately, estimates of $\eta_{l}$ (total within-lanugage variability), $\sigma_{l}$ (model-side variability), and $\tau_{l}$ (test data variability) provide the researcher with estimates of what is driving the variation in their data. High values of $\sigma_{l}^{2}$ (model-side variability) suggest that the learned distribution over responses may have high entropy or, in the case of finetuning, that the architecture may be highly sensitive to random shuffling of batches due to e.g. small amounts of training data. On the other hand, high values of $\tau_{l}^{2}$ (test data variability) and $\nu_{m}^{2}$ (between-language variance) indicate that a model’s performance is particularly sensitive to the exact composition of a test set or benchmark and may suggest lower ability to generalize due to overfitting. Variance components are also intimately tied to statistical significance: we can, roughly speaking, judge an estimate as being statistically significant if it lies more than two standard errors from zero (e.g. Gelman, 2018).
Estimating model- and data-side SD
@@
To estimate the boot-to-boot variability $\tau_{l}$, we first compute the F1 score variance over the $B$ bootstrap datasets separately for each of the $S$ seeds.We then construct the average estimator $\bar{\tau}_{l}=\sum_{s=1}^{S}\hat{\tau}_{ls}/S$. The standard error of $\hat{\tau}_{l}$ is estimated using the usual sample standard deviation formula over seeds.This gives us the plug-in estimate of the standard error of the averaged estimator $\hat{\text{se}}(\bar{\tau}_{l})=\hat{\text{se}}(\hat{\tau}_{ls})/\sqrt{S}$.
@@
In most leaderboard-style scenarios, aggregated scalar performance measures such as the arithmetic mean $\bar{x}^{(m)}_{1:L}=\sum_{l}x^{(m)}_{l}/L$ tend to be used instead of the full score vectors. Its popularity of the arithmetic mean can be explained by how simple it is to compute, as well as the closed-form expression for its standard deviation $\text{sd}(x^{(m)}_{1},\dots,x^{(m)_{L}})/\sqrt{L}$ using only the between-language SD $\nu_{m}$ and the within-language SDs $\eta_{ml}$. Other aggregation functions, such as the geometric mean, or median may also be used, although they may not be as well-behaved in terms of standard error. For such aggregation functions, closed form expressions of the SD do not typically exist and must be estimated using resampling.
4 Task 1: LLM-based question answering
As our first case study, we focus on multilingual question answering and evaluate four LLMs on all subsets of using XQUAD (Artetxe et al., 2020). Specifically, we use aya-expanse-8B (Dang et al., 2024), TowerInstruct-Mistral-7B-v0.2 (Alves et al., 2024), Google’s gemma2-9b and finally Clarus-7B-v0.3. We evaluate using token-level F1 score, using the standard lm-evaluation-harness implementation. Details of the corpus, experimental settings and hardware used are in the appendix.
@@ 4.1 Variance components
The lm-eval-harness library we use to run our experiments automatically computes bootstrapped standard errors for F1. In addition to the bootstrap SE, we incorporate model-side uncertainty by sampling 5 answers for each question from each model. The summarized variance components are displayed in Table 1. Full decomposition is available in Table 10 in the appendix.
@@
Clarus-7B 8.92 - - -
TowerInstruct-7B 11.03 - - -
aya-expanse-8B 12.01 - - -
gemma2-9B 5.05 - - -
@@
Clarus-7B 0.89 0.41 0.40 1.67
TowerInstruct-7B 0.73 0.32 0.19 1.21
aya-expanse-8B 1.47 0.18 1.18 1.73
gemma2-9B 1.21 0.16 0.93 1.51
@@
Clarus-7B 0.70 0.29 0.24 1.27
TowerInstruct-7B 0.59 0.28 0.15 1.05
aya-expanse-8B 1.24 0.11 1.06 1.42
gemma2-9B 0.95 0.10 0.80 1.18
@@
Clarus-7B 0.53 0.33 0.16 1.25
TowerInstruct-7B 0.41 0.18 0.12 0.74
aya-expanse-8B 0.76 0.30 0.25 1.23
gemma2-9B 0.74 0.20 0.48 1.09
@@
Clarus-7B TowerInstruct-7B aya-expanse-8b
TowerInstruct-7B aya-expanse-8b gemma-2-9b aya-expanse-8b gemma-2-9b gemma-2-9b

### v2

# Beyond statistical significance: Quantifying uncertainty and statistical variability in multilingual and multitask NLP evaluation

@@ Abstract
We introduce a set of resampling-based methods for quantifying uncertainty and statistical precision of evaluation metrics in multilingual and/or multitask NLP benchmarks. We show how experimental variation in performance scores arises from both model and data-related sources, and that accounting for both of them is necessary to avoid substantially underestimating the overall variability over hypothetical replications. Using multilingual question answering, machine translation, and named entity recognition as example tasks, we also demonstrate how resampling methods are useful for quantifying the replication uncertainty of various quantities used in leaderboards such as model rankings and pairwise differences between models.
Beyond statistical significance: Quantifying uncertainty and statistical variability in multilingual and multitask NLP evaluation
@@ 1 Introduction
Over the last several years, multilingual research has undergone exponential growth within NLP, both in terms of model capabilities as well as evaluation datasets. Since no paper can evaluate on all languages, it is important to determine to what extent research findings generalize from one language and from one task to another. This is particularly true in recent years with the proliferation of large language models (LLMs), where the goal often is to draw conclusions about the models that are not task or language-dependent.
Evaluation paradigms
Despite this need, evaluation setups in multilingual NLP research tend to focus on within-language evaluation and fall into two broad categories. In the first category, the same experiments are repeated across several languages and the results analyzed separately. The inferences are typically reported as one collection, e.g. “Model A significantly outperforms model B on English–Finnish and English–Turkish but not English–German” or “Model A outperforms model B on 13 out of 43 language pairs.” The second variant, depicted in Figure 1, is the “leaderboard-type” scenario in which performance is evaluated on each of the $L$ languages after which the numbers are distilled into a single performance measure using a pre-specified aggregation function such as the arithmetic mean. Given the aggregate scores, the main interest may then be to rank the $M$ models based on the single score or to compare differences between observed scores.
@@
Statistical significance and effect sizes
[…]
To estimate such effect sizes requires specifying the source(s) of random variation in our experiments. Typically, this variation has been treated as arising from one of two sources: data-side variability and model-side variability which we describe below.
Data-side variability
The former quantifies how differently each model would perform if a slightly different test set were used for a given task. In practice this is typically done using resampled versions of the original test set, for example using the bootstrap resampling algorithm (Efron, 1979).
@@
From a sampling perspective, data-side variability can be interpreted as sampling error associated with the process of constructing a test set for a given task by sampling data from a larger set of all possible test sets. Another source of data-related variability that appears in leaderboards and other multi-task evaluations is the choice of evaluation tasks. When ranking is performed based on aggregated quantities, the overall performance and rankings of models will vary as a function of which tasks are chosen. This can be understood as the sampling error associated with creating an evaluation benchmark out of the larger population of all possible ones that could have been constructed.
Model-side variability
@@
There is also model-side variability that will be present even if the data set is held constant. With LLMs, decoding often involves random sampling of tokens, especially with generative tasks like question answering. By drawing multiple responses for each question, the performance will fluctuate randomly around some average. Even if the inference algorithm is deterministic (e.g. beam search), any potential finetuning will likely be nondeterministic due to, for example, random shuffling of minibatches and weight initialization. This will yield slightly different parameters from run to run, which will potentially yield different responses and different values of the performance metric. From a sampling perspective, model-side uncertainty can be seen as drawing samples from the probability distribution over possible responses defined by the model.
Summary
@@
In this paper, we provide a new perspective on how resampling-based methods can be useful in analysis of experimental results in multilingual and multitask NLP.
@@
Through connections to statistics, we show how experimental variance can be decomposed into between and within-task components ($\nu$ and $\eta_{l}$) and how the latter further decomposes into model and data-based components ($\sigma_{l}$, $\tau_{l}$) which arise naturally from the structure of the experimental data.
@@
Using question answering (QA), machine translation (MT), and named entity recognition (NER) as case studies, our results demonstrate that none of these sources of variation are negligible. This suggests that only analyzing one source of error may underestimate the total variation and expose researchers to the risk of drawing incorrect conclusions about the relative merits of models.
@@
We show how resampling and estimates of between and within-language variance components can be used to derive uncertainty-aware estimates of more complex quantities such as relative rankings of models and approximate distributions of pairwise performance differences between them.
@@
All of our methods run in seconds to minutes and present no major computational bottlenecks on top of the overall inference time complexity. We provide an implementation of our approach in a toolkit which we make freely available at https://github.com/j0ma/reuben. [^1] The name reuben is an acronym for “REsampling-based Uncertainty Bounds for Evaluating NLP.
@@ 2 Related work
In contrast, Ulmer et al. (2022) advocate for understanding model performance variation based on multiple model training runs instead of resampling, and using this model-side variation to assess statistical significance. There is no clear consensus though, and others (e.g. Bethard, 2022) expressly warn against using random seeds to provide estimates of score distributions based on the argument that random seeds should be optimized like other hyperparameters. Some work in NLP has also evaluated different sources of variation in evaluation of MT metrics (Xiang et al., 2022).
@@
There has also been a healthy debate regarding best practices in test set construction and more generally of evaluation using held-out datasets. Gorman and Bedrick (2019) and others (e.g. Kodner et al., 2023; Liu and Dorr, 2024) argue that we should randomize our splits and evaluate on several versions of our test sets to reduce the chance of false positives. Søgaard et al. (2021) instead favor using approaches based on adversarial splitting, arguing that naive randomization will underestimate the variation.
Benchmarks have received much attention in the general machine learning community. An excellent survey is Dehghani et al. (2021) where the authors discuss the lifecycle of overall ML benchmarks, how they become stale as well as what is random when models are evaluated. The question of what aggregation function to use in leaderboards is also explored by Tatiana and Valentin (2021) where the authors explore using both the arithmetic, geometric and harmonic means. In particular, they show that results may be wildly different depending on what aggregation measure is used. Longjohn et al. (2025) also survey and develop resampling-based and Bayesian methods for popular computer vision benchmarks for LLMs.
@@
There exists a substantial literature (e.g. Nikitin et al., 2024; Da et al., 2025; Chen et al., 2024, 2025; Yang et al., 2025; Ye et al., 2024; Wagner et al., 2024; Blackwell et al., 2025) on a task referred to as uncertainty quantification in NLP, where the task is to estimate some properties of an LLM’s distribution over continuations, $p(x|x_{\text{prompt}})$, that predicts whether the model is able to produce the correct answer to this question. While this task quantifies per-example uncertainty, our efforts instead focus on experimental variation. i.e. estimating the noise we would expect to see under repeated experiments and hypothetical evaluation suites.
@@ 3 Statistical background
Under infinite replications, our observed values $x^{(m)}_{l}$ values will center around some mean value and exhibit a degree of fluctuation around it, representing measurement error arising from both model-side randomness (e.g. nondeterministic decoding) as well as sampling error of the test set. We can represent this as $x^{(m)}_{lr}=\mu^{(m)}_{l}+\varepsilon^{\text{repl}}_{lr}$ where $\varepsilon^{\text{repl}}_{lr}$ represents the deviation of observed scores from the language-specific mean $\mu^{(m)}_{l}$. In addition to this “within-language” variation arising from replication, there is also “between-language” variability related to what tasks are included in the leaderboard. Concretely, each $\mu^{(m)}_{l}$ also differs from the global mean performance across all tasks by some amount, i.e. $\mu^{(m)}_{l}=\bar{\mu}^{(m)}+\varepsilon^{\text{lang}}_{l}$. This lets us decompose the overall observations into $x^{(m)}_{lr}=\bar{\mu}^{(m)}+\varepsilon^{\text{lang}}_{l}+\varepsilon^{\text{repl}}_{lr}$. Taking the variance of both sides, we obtain a first-principles variance decomposition into between and within-language components $\mathbb{V}[x^{(m)}_{lr}]=\mathbb{V}[\varepsilon^{\text{lang}}_{l}]+\mathbb{V}[\varepsilon^{\text{repl}}_{lr}]=\nu_{m}^{2}+\eta^{2}_{lm}$ where $\nu_{m}^{2}$ and $\eta^{2}_{lm}$ refer to between and within-language variation of model $m$, respectively.
@@
In our analysis, we use both seed and bootstrap-based resampling to yield a total of $R=SB$ replications per language. While the $B$ resamples may be much cheaper to obtain than the $S$ model re-instantiations in terms of time cost, we feel that using both is helpful to properly estimate the decomposition as well as to understand whether the sources contribute equally to the total within-language variance for each language.
@@
Most immediately, estimates of $\eta_{l}$ (total within-language variability), $\sigma_{l}$ (model-side variability), and $\tau_{l}$ (test data variability) provide the researcher with estimates of what is driving the variation in their data. High values of $\sigma_{l}^{2}$ (model-side variability) suggest that the learned distribution over responses may have high entropy or, in the case of finetuning, that the architecture may be highly sensitive to random shuffling of batches due to e.g. small amounts of training data. On the other hand, high values of $\tau_{l}^{2}$ (test data variability) and $\nu_{m}^{2}$ (between-language variance) indicate that a model’s performance is particularly sensitive to the exact composition of a test set or benchmark and may suggest lower ability to generalize due to overfitting. Variance components are also tied to statistical significance: we can, roughly speaking, judge an estimate as being statistically significant if it lies more than two standard errors from zero (e.g. Gelman, 2018).
Estimating model and data-side SD
@@
To estimate the boot-to-boot variability $\tau_{l}$, we first compute the F1 score variance over the $B$ bootstrap datasets separately for each of the $S$ seeds.
We then construct the average estimator $\bar{\tau}_{l}=\sum_{s=1}^{S}\hat{\tau}_{ls}/S$. The standard error of $\hat{\tau}_{l}$ is estimated using the usual sample standard deviation formula over seeds. This gives us the plug-in estimate of the standard error of the averaged estimator $\hat{\text{se}}(\bar{\tau}_{l})=\hat{\text{se}}(\hat{\tau}_{ls})/\sqrt{S}$.
@@
In most leaderboard-style scenarios, aggregated scalar performance measures such as the arithmetic mean $\bar{x}^{(m)}_{1:L}=\sum_{l}x^{(m)}_{l}/L$ tend to be used instead of the full score vectors. Its popularity of the arithmetic mean can be explained by how simple it is to compute, as well as the closed-form expression for its standard deviation $\text{sd}(x^{(m)}_{1},\dots,x^{(m)_{L}})/\sqrt{L}$ using only the between-language SD $\nu_{m}$ and the within-language SDs $\eta_{ml}$. Other aggregation functions, such as the geometric mean, or median may also be used, although they may not be as well-behaved in terms of standard error. For such aggregation functions, the SD typically does not exist in closed form and must be estimated using resampling.
4 Task 1: Question answering with LLMs
As our first case study, we focus on multilingual question answering and evaluate four LLMs on all subsets of XQuAD (Artetxe et al., 2020). Specifically, we use aya-expanse-8B (Dang et al., 2024), TowerInstruct-Mistral-7B-v0.2 (Alves et al., 2024), Google’s gemma2-9b and finally Clarus-7B-v0.3. We evaluate using token-level F1 score, using the standard lm-evaluation-harness implementation. Details of the corpus, experimental settings and hardware used are in Section B.1 in the Appendix.
@@ 4.1 Variance components
The lm-evaluation-harness library we use to run our experiments automatically computes bootstrapped standard errors for F1. In addition to the bootstrap SE, we incorporate model-side uncertainty by sampling 5 answers for each question from each model. The summarized variance components are displayed in Table 1. The full decomposition is available in Table 10 in the Appendix.
@@
Clarus 7B 8.92 - - -
TowerInstruct 7B 11.03 - - -
Aya Expanse 8B 12.01 - - -
Gemma 2 9B 5.05 - - -
@@
Clarus 7B 0.89 0.41 0.40 1.67
TowerInstruct 7B 0.73 0.32 0.19 1.21
Aya Expanse 8B 1.47 0.18 1.18 1.73
Gemma 2 9B 1.21 0.16 0.93 1.51
@@
Clarus 7B 0.70 0.29 0.24 1.27
TowerInstruct 7B 0.59 0.28 0.15 1.05
Aya Expanse 8B 1.24 0.11 1.06 1.42
Gemma 2 9B 0.95 0.10 0.80 1.18
@@
Clarus 7B 0.53 0.33 0.16 1.25
TowerInstruct 7B 0.41 0.18 0.12 0.74
Aya Expanse 8B 0.76 0.30 0.25 1.23
Gemma 2 9B 0.74 0.20 0.48 1.09
@@
Clarus 7B TowerInstruct 7B Aya Exp. 8B
TowerInstruct 7B Aya Exp. 8B Gemma 2 9B Aya Exp. 8B Gemma 2 9B Gemma 2 9B